OpenStudy (anonymous):

The function t(x) = 3x + 5 determines how many cans of corn kernels a food truck needs to stock, where x is the number of shifts the crew is going to work in the truck. The crew uses c(t(x)) to find the amount of money to spend on corn. The function c(x) = 2x + 3. Solve for how much money must be spent when the crew is going to work 3 shifts. 11 14 31 43

OpenStudy (anonymous):

Let us first find c(t(x)) we have c(x) = 2x + 3 Now substituting x= t(x) in c(x) we find: c(t(x)) = 2(3x + 5) + 3 = 6x +10 + 3= 6x+ 13 Next let us substitute x= 3 as x is the number of shifts and crew is going to work 3 shifts Thus we find c(t(3)) = 6(3)+ 13 = 18 + 13 = 31 hence amount of money to be spent is 31 Hence third option is the correct answer..

OpenStudy (anonymous):

Can you help me on one more probelm @dpasingh

OpenStudy (anonymous):

@gabylovesu Sure if I would be able to do them

OpenStudy (anonymous):

Ok!

OpenStudy (anonymous):

Pleae post as new question in ask question section by mentioning me